### 29 Mar 2026 v2

# On the Classification of Two-Dimensional Algebras

Abstract. We provide a clarification of the classification of two-dimensional algebras over an arbitrary base field. Using this clarification, we determine the number of non-isomorphic two-dimensional algebras over a finite field.
1. Introduction
The number of non-isomorphic two-dimensional algebras over a finite field $\mathbb{F}$ with $|\mathbb{F}|=q$ was computed in [1] as
$q^{4}+q^{3}+4q^{2}+4q+7$, if $char(\mathbb{F})\neq 2,3$,
$q^{4}+q^{3}+4q^{2}+3q+6$, if $char(\mathbb{F})=2$,
$q^{4}+q^{3}+4q^{2}+4q+6$, if $char(\mathbb{F})=3$.
For another approach to calculating this number, see [2].
Our original aim was to derive this number using the classification result [3]. However, the difference between the above formula and the number obtained from the classification result forced us to recheck the classification one more time. It turns out that some corrections are needed in the classification of the fifth family of two-dimensional algebras in [3].
In this paper, we first provide these corrections and then use the revised classification to determine the number of non-isomorphic two-dimensional algebras over a finite field.
The classification result presented in [3] is the following.
Theorem $1.$
Any non-trivial two-dimensional algebra over a field $\mathbb{F}$ with $char(\mathbb{F})\neq 2,3$ is isomorphic to exactly one of the following algebras, given by their matrices of structure constants.
  • $A_{1}(\mathbf{c})=\left(\begin{array}[]{cccc}\alpha_{1}&\alpha_{2}&1+\alpha_{2}&\alpha_{4}\\ \beta_{1}&-\alpha_{1}&1-\alpha_{1}&-\alpha_{2}\end{array}\right),\ \mbox{where}\ \mathbf{c}=(\alpha_{1},\alpha_{2},\alpha_{4},\beta_{1})\in\mathbb{F}^{4},$
  • $A_{2}(\mathbf{c})=\left(\begin{array}[]{cccc}\alpha_{1}&0&0&\alpha_{4}\\ 1&\beta_{2}&1-\alpha_{1}&0\end{array}\right),\ \mbox{where}\ \mathbf{c}=(\alpha_{1},\alpha_{4},\beta_{2})\in\mathbb{F}^{3},\ \alpha_{4}\neq 0,$
  • $A_{3}(\mathbf{c})=\left(\begin{array}[]{cccc}\alpha_{1}&0&0&\alpha_{4}\\ 0&\beta_{2}&1-\alpha_{1}&0\end{array}\right)\simeq\left(\begin{array}[]{cccc}\alpha_{1}&0&0&a^{2}\alpha_{4}\\ 0&\beta_{2}&1-\alpha_{1}&0\end{array}\right),$ where $\mathbf{c}=(\alpha_{1},\alpha_{4},\beta_{2})\in\mathbb{F}^{3}$ and $0\neq a\in\mathbb{F}$,
  • $A_{4}(\mathbf{c})=\left(\begin{array}[]{cccc}0&1&1&0\\ \beta_{1}&\beta_{2}&1&-1\end{array}\right),\ \mbox{where}\ \mathbf{c}=(\beta_{1},\beta_{2})\in\mathbb{F}^{2},$
  • $A_{5}(\mathbf{c})=\left(\begin{array}[]{cccc}\alpha_{1}&0&0&0\\ 1&2\alpha_{1}-1&1-\alpha_{1}&0\end{array}\right),\ \mbox{where}\ \mathbf{c}=\alpha_{1}\in\mathbb{F},$
  • $A_{6}(\mathbf{c})=\left(\begin{array}[]{cccc}\alpha_{1}&0&0&\alpha_{4}\\ 1&1-\alpha_{1}&-\alpha_{1}&0\end{array}\right),\ \mbox{where}\ \mathbf{c}=(\alpha_{1},\alpha_{4})\in\mathbb{F}^{2},\ \alpha_{4}\neq 0,$
  • $A_{7}(\mathbf{c})=\left(\begin{array}[]{cccc}\alpha_{1}&0&0&\alpha_{4}\\ 0&1-\alpha_{1}&-\alpha_{1}&0\end{array}\right)\simeq\left(\begin{array}[]{cccc}\alpha_{1}&0&0&a^{2}\alpha_{4}\\ 0&1-\alpha_{1}&-\alpha_{1}&0\end{array}\right),$ where $\mathbf{c}=(\alpha_{1},\alpha_{4})\in\mathbb{F}^{2}$ and $0\neq a\in\mathbb{F}$,
  • $A_{8}(\mathbf{c})=\left(\begin{array}[]{cccc}0&1&1&0\\ \beta_{1}&1&0&-1\end{array}\right),\ \mbox{where}\ \mathbf{c}=\beta_{1}\in\mathbb{F},$
  • $A_{9}(\mathbf{c})=\left(\begin{array}[]{cccc}\frac{1}{3}&0&0&0\\ 1&\frac{2}{3}&-\frac{1}{3}&0\end{array}\right),$
  • $A_{10}(\mathbf{c})=\left(\begin{array}[]{cccc}0&1&1&1\\ \beta_{1}&0&0&-1\end{array}\right)\simeq\left(\begin{array}[]{cccc}0&1&1&1\\ \beta^{\prime}_{1}(a)&0&0&-1\end{array}\right),$ where polynomial $(\beta_{1}t^{3}-3t-1)(\beta_{1}t^{2}+\beta_{1}t+1)(\beta_{1}^{2}t^{3}+6\beta_{1}t^{2}+3\beta_{1}t+\beta_{1}-2)$ has no root in $\mathbb{F}$, $a\in\mathbb{F}$ and $\beta^{\prime}_{1}(t)=\frac{(\beta_{1}^{2}t^{3}+6\beta_{1}t^{2}+3\beta_{1}t+\beta_{1}-2)^{2}}{(\beta_{1}t^{2}+\beta_{1}t+1)^{3}}$,
  • $A_{11}(\mathbf{c})=\left(\begin{array}[]{cccc}0&0&0&1\\ \beta_{1}&0&0&0\end{array}\right)\simeq\left(\begin{array}[]{cccc}0&0&0&1\\ a^{3}\beta_{1}^{\pm 1}&0&0&0\end{array}\right),$ where polynomial $\beta_{1}-t^{3}$ has no root in $\mathbb{F}$, $\mathbf{c}=\beta_{1}\neq 0$ and $0\neq a\in\mathbb{F}$,
  • $A_{12}(\mathbf{c})=\left(\begin{array}[]{cccc}0&1&1&0\\ \beta_{1}&0&0&-1\end{array}\right)\simeq\left(\begin{array}[]{cccc}0&1&1&0\\ a^{2}\beta_{1}&0&0&-1\end{array}\right),\ \mbox{where}\ \mathbf{c}=\beta_{1}\in\mathbb{F},\ 0\neq a\in\mathbb{F},$
  • $A_{13}=\left(\begin{array}[]{cccc}0&0&0&0\\ 1&0&0&0\end{array}\right).$
Any non-trivial 2-dimensional algebra over a field $\mathbb{F}$, $char.(\mathbb{F})=2$, is isomorphic to only one of the following listed, by their matrices of structure constants, algebras. Moreover, their automorphism groups and derivations, in that basis, are as follows:
  • $A_{1,2}(\mathbf{c})=\left(\begin{array}[]{cccc}\alpha_{1}&\alpha_{2}&\alpha_{2}+1&\alpha_{4}\\ \beta_{1}&\alpha_{1}&1+\alpha_{1}&\alpha_{2}\end{array}\right),\ \mbox{where}\ \mathbf{c}=(\alpha_{1},\alpha_{2},\alpha_{4},\beta_{1})\in\mathbb{F}^{4},$
  • $A_{2,2}(\mathbf{c})=\left(\begin{array}[]{cccc}\alpha_{1}&0&0&\alpha_{4}\\ 1&\beta_{2}&1+\alpha_{1}&0\end{array}\right),\ \mbox{where}\ \mathbf{c}=(\alpha_{1},\alpha_{4},\beta_{2})\in\mathbb{F}^{3},\ \alpha_{4}\neq 0,$
    $A_{2,2}(\alpha_{1},0,1)=\left(\begin{array}[]{cccc}\alpha_{1}&0&0&0\\ 1&1&1+\alpha_{1}&0\end{array}\right),\ \mbox{where}\ \alpha_{1}\in\mathbb{F},$
  • $A_{3,2}(\mathbf{c})=\left(\begin{array}[]{cccc}\alpha_{1}&0&0&\alpha_{4}\\ 0&\beta_{2}&1+\alpha_{1}&0\end{array}\right)\simeq\left(\begin{array}[]{cccc}\alpha_{1}&0&0&a^{2}\alpha_{4}\\ 0&\beta_{2}&1+\alpha_{1}&0\end{array}\right),$ where $\mathbf{c}=(\alpha_{1},\alpha_{4},\beta_{2})\in\mathbb{F}^{3}$ and $0\neq a\in\mathbb{F}$,
  • $A_{4,2}(\mathbf{c})=\left(\begin{array}[]{cccc}\alpha_{1}&1&1&0\\ \beta_{1}&\beta_{2}&1+\alpha_{1}&1\end{array}\right)\simeq\left(\begin{array}[]{cccc}\alpha_{1}&1&1&0\\ \beta_{1}+(1+\beta_{2})a+a^{2}&\beta_{2}&1+\alpha_{1}&1\end{array}\right),\ \mbox{where}\ \mathbf{c}=(\alpha_{1},\beta_{1},\beta_{2})\in\mathbb{F}^{2},$
  • $A_{5,2}(\mathbf{c})=\left(\begin{array}[]{cccc}\alpha_{1}&0&0&\alpha_{4}\\ 1&1+\alpha_{1}&\alpha_{1}&0\end{array}\right),\ \mbox{where}\ \mathbf{c}=(\alpha_{1},\alpha_{4})\in\mathbb{F}^{2},\ \alpha_{4}\neq 0,\\ \ A_{5,2}(1,0)=\left(\begin{array}[]{cccc}1&0&0&0\\ 1&0&1&0\end{array}\right),$
  • $A_{6,2}(\mathbf{c})=\left(\begin{array}[]{cccc}\alpha_{1}&0&0&\alpha_{4}\\ 0&1+\alpha_{1}&\alpha_{1}&0\end{array}\right)\simeq\left(\begin{array}[]{cccc}\alpha_{1}&0&0&a^{2}\alpha_{4}\\ 0&1+\alpha_{1}&\alpha_{1}&0\end{array}\right),$ where $\mathbf{c}=(\alpha_{1},\alpha_{4})\in\mathbb{F}^{2}$ and $0\neq a\in\mathbb{F}$,
  • $A_{7,2}(\mathbf{c})=\left(\begin{array}[]{cccc}\alpha_{1}&1&1&0\\ \beta_{1}&1+\alpha_{1}&\alpha_{1}&1\end{array}\right)\simeq\left(\begin{array}[]{cccc}\alpha_{1}&1&1&0\\ \beta_{1}+a\alpha_{1}+a+a^{2}&1+\alpha_{1}&\alpha_{1}&1\end{array}\right),\\ \mbox{where}\ \mathbf{c}=(\alpha_{1},\beta_{1})\in\mathbb{F}^{2},\ a\in\mathbb{F},$
  • $A_{8,2}(\mathbf{c})=\left(\begin{array}[]{cccc}0&1&1&1\\ \beta_{1}&0&0&1\end{array}\right)\simeq\left(\begin{array}[]{cccc}0&1&1&1\\ \beta^{\prime}_{1}(a)&0&0&1\end{array}\right),$ where polynomial $(\beta_{1}t^{3}+t+1)(\beta_{1}t^{2}+\beta_{1}t+1)$ has no root in $\mathbb{F}$, $a\in\mathbb{F}$ and $\beta^{\prime}_{1}(t)=\frac{(\beta_{1}^{2}t^{3}+\beta_{1}t+\beta_{1})^{2}}{(\beta_{1}t^{2}+\beta_{1}t+1)^{3}}$,
  • $A_{9,2}(\mathbf{c})=\left(\begin{array}[]{cccc}0&0&0&1\\ \beta_{1}&0&0&0\end{array}\right)\simeq\left(\begin{array}[]{cccc}0&0&0&1\\ a^{3}\beta_{1}^{\pm 1}&0&0&0\end{array}\right),\ \mbox{where}\ \mathbf{c}=\beta_{1}\in\mathbb{F},\\ \ 0\neq a\in\mathbb{F},$ polynomial $\beta_{1}+t^{3}$ has no root in $\mathbb{F}$,
  • $A_{10,2}(\mathbf{c})=\left(\begin{array}[]{cccc}1&1&1&0\\ \beta_{1}&1&1&1\end{array}\right)\simeq\left(\begin{array}[]{cccc}1&1&1&0\\ \beta_{1}+a+a^{2}&1&1&1\end{array}\right),\ \mbox{where}\ \mathbf{c}=\beta_{1}\in\mathbb{F},\ a\in\mathbb{F},$
  • $A_{11,2}=\left(\begin{array}[]{cccc}0&1&1&0\\ \beta_{1}&0&0&1\end{array}\right)\simeq\left(\begin{array}[]{cccc}0&1&1&0\\ b^{2}(\beta_{1}+a^{2})&0&0&1\end{array}\right),$ where $0\neq b$,
    $a,b\in\mathbb{F}$,
  • $A_{12,2}=\left(\begin{array}[]{cccc}0&0&0&0\\ 1&0&0&0\end{array}\right).$
Any non-trivial 2-dimensional algebra over a field $\mathbb{F}$, $char.(\mathbb{F})=3$, is isomorphic to only one of the following listed, by their matrices of structure constants, algebras. Moreover, their automorphism groups and derivations, in that basis, are as follows:
  • $A_{1,3}(\mathbf{c})=\left(\begin{array}[]{cccc}\alpha_{1}&\alpha_{2}&\alpha_{2}+1&\alpha_{4}\\ \beta_{1}&-\alpha_{1}&1-\alpha_{1}&-\alpha_{2}\end{array}\right),\ \mbox{where}\ \mathbf{c}=(\alpha_{1},\alpha_{2},\alpha_{4},\beta_{1})\in\mathbb{F}^{4},$
  • $A_{2,3}(\mathbf{c})=\left(\begin{array}[]{cccc}\alpha_{1}&0&0&\alpha_{4}\\ 1&\beta_{2}&1-\alpha_{1}&0\end{array}\right),\ \mbox{where}\ \mathbf{c}=(\alpha_{1},\alpha_{4},\beta_{2})\in\mathbb{F}^{3},\ \alpha_{4}\neq 0,$
  • $A_{3,3}(\mathbf{c})=\left(\begin{array}[]{cccc}\alpha_{1}&0&0&\alpha_{4}\\ 0&\beta_{2}&1-\alpha_{1}&0\end{array}\right)\simeq\left(\begin{array}[]{cccc}\alpha_{1}&0&0&a^{2}\alpha_{4}\\ 0&\beta_{2}&1-\alpha_{1}&0\end{array}\right),$ where $\mathbf{c}=(\alpha_{1},\alpha_{4},\beta_{2})\in\mathbb{F}^{3}$ and $0\neq a\in\mathbb{F}$,
  • $A_{4,3}(\mathbf{c})=\left(\begin{array}[]{cccc}0&1&1&0\\ \beta_{1}&\beta_{2}&1&-1\end{array}\right),\ \mbox{where}\ \mathbf{c}=(\beta_{1},\beta_{2})\in\mathbb{F}^{2},$
  • $A_{5,3}(\mathbf{c})=\left(\begin{array}[]{cccc}\alpha_{1}&0&0&0\\ 1&2\alpha_{1}-1&1-\alpha_{1}&0\end{array}\right),\ \mbox{where}\ \mathbf{c}=\alpha_{1}\in\mathbb{F},$
  • $A_{6,3}(\mathbf{c})=\left(\begin{array}[]{cccc}\alpha_{1}&0&0&\alpha_{4}\\ 1&1-\alpha_{1}&-\alpha_{1}&0\end{array}\right),\ \mbox{where}\ \mathbf{c}=(\alpha_{1},\alpha_{4})\in\mathbb{F}^{2},\ \alpha_{4}\neq 0,$
  • $A_{7,3}(\mathbf{c})=\left(\begin{array}[]{cccc}\alpha_{1}&0&0&\alpha_{4}\\ 0&1-\alpha_{1}&-\alpha_{1}&0\end{array}\right)\simeq\left(\begin{array}[]{cccc}\alpha_{1}&0&0&a^{2}\alpha_{4}\\ 0&1-\alpha_{1}&-\alpha_{1}&0\end{array}\right),$ where $\mathbf{c}=(\alpha_{1},\alpha_{4})\in\mathbb{F}^{2}$ and $0\neq a\in\mathbb{F}$,
  • $A_{8,3}(\mathbf{c})=\left(\begin{array}[]{cccc}0&1&1&0\\ \beta_{1}&1&0&-1\end{array}\right),\ \mbox{where}\ \mathbf{c}=\beta_{1}\in\mathbb{F},$
  • $A_{9,3}(\beta_{1})=\left(\begin{array}[]{cccc}0&1&1&1\\ \beta_{1}&0&0&-1\end{array}\right)\simeq\left(\begin{array}[]{cccc}0&1&1&1\\ \beta^{\prime}_{1}(a)&0&0&-1\end{array}\right),$ where polynomial
    $(\beta_{1}-t^{3})(\beta_{1}t^{2}+\beta_{1}t+1)(\beta_{1}^{2}t^{3}+\beta_{1}-2)$ has no root in $\mathbb{F}$, $a\in\mathbb{F}$ and $\beta^{\prime}_{1}(t)=\frac{(\beta_{1}^{2}t^{3}+\beta_{1}-2)^{2}}{(\beta_{1}t^{2}+\beta_{1}t+1)^{3}}$. There is no such algebra as far as (3,q1)=1(3,q-1)=1 every element has a cubic root, $\beta_{1}=t^{3}$ has a solution.
  • $A_{10,3}(\beta_{1})=\left(\begin{array}[]{cccc}0&0&0&1\\ \beta_{1}&0&0&0\end{array}\right)\simeq\left(\begin{array}[]{cccc}0&0&0&1\\ a^{3}\beta_{1}^{\pm 1}&0&0&0\end{array}\right),$ where polynomial
    $\beta_{1}-t^{3}$ has no root and $0\neq a\in\mathbb{F}$,
  • $A_{11,3}(\beta_{1})=\left(\begin{array}[]{cccc}0&1&1&0\\ \beta_{1}&0&0&-1\end{array}\right)\simeq\left(\begin{array}[]{cccc}0&1&1&0\\ a^{2}\beta_{1}&0&0&-1\end{array}\right),\ \mbox{where}\ \beta_{1}\in\mathbb{F},\\ \ 0\neq a\in\mathbb{F},$
  • $A_{12,3}=\left(\begin{array}[]{cccc}1&0&0&0\\ 1&-1&-1&0\end{array}\right),$
  • $A_{13,3}=\left(\begin{array}[]{cccc}0&0&0&0\\ 1&0&0&0\end{array}\right).$
Remark 1.2.
Note that any two algebras taken from different classes listed above are mutually non-isomorphic. This follows from the classification method presented in [3]. It can also be justified using distinguishing invariants. A comparison of their automorphism groups and derivation algebras given in [4] reduces the problem to only a few pairs of classes. For example, in $char.(\mathbb{F})\neq 2,3$ case the corresponding pairs are $(A_{2},A_{4})$, $(A_{6},A_{8})$ and all mutual pairs of $A_{10}$, $A_{11}$ and $A_{12}$. The idempotent equation for $A_{2}$ ($A_{3}$) is reduced to a $3^{rd}$-order(respectively, $2^{nd}$-order) polynomial equation, therefore they are not isomorphic. The same argument works in $A_{6}$ with $A_{8}$ case as well. Algebras $A_{10}$, $A_{11}$ and $A_{12}$ are commutative, 33 power associative and for $x=ue_{1}+ve_{2}$ one has, respectively, $x^{3}=x(u^{2}\beta_{1}+uv\beta_{1}+v^{2})$, $x^{3}=xuv\beta_{1}$, $x^{3}=x(u^{2}\beta_{1}+v^{2})$. This implies that these algebras are mutually non-isomorphic. Similar approach works in other characteristics as well.
2. On the Classification of Two-Dimensional Algebras over an Arbitrary Base Field
In this section we correct certain inaccuracies in the classification result presented in [3].
Recall that a two-dimensional algebra given by the matrix
$A=\left(\begin{array}[]{cccc}\alpha_{1}&\alpha_{2}&\alpha_{3}&\alpha_{4}\\ \beta_{1}&\beta_{2}&\beta_{3}&\beta_{4}\end{array}\right)$
means the algebra with multiplication
$e_{1}e_{1}=\alpha_{1}e_{1}+\beta_{1}e_{2},\ e_{1}e_{2}=\alpha_{2}e_{1}+\beta_{2}e_{2},\ e_{2}e_{1}=\alpha_{3}e_{1}+\beta_{3}e_{2},\ e_{2}e_{2}=\alpha_{4}e_{1}+\beta_{4}e_{2}.$
Our corrections concern only the following case considered in that paper, and we follow the notation used there.
The fifth subset in the $char(\mathbb{F})\neq 2,3$ case. In this case
$A=\left(\begin{array}[]{cccc}\alpha_{1}&\alpha_{2}&\alpha_{2}&\alpha_{4}\\ \beta_{1}&-\alpha_{1}&-\alpha_{1}&-\alpha_{2}\end{array}\right)\ \mbox{and if}\ g^{-1}=\left(\begin{array}[]{cc}\xi_{1}&\eta_{1}\\ \xi_{2}&\eta_{2}\end{array}\right),\ \Delta=\xi_{1}\eta_{2}-\xi_{2}\eta_{1}\ \mbox{and}$
$A^{\prime}=\left(\begin{array}[]{cccc}\alpha^{\prime}_{1}&\alpha^{\prime}_{2}&\alpha^{\prime}_{2}&\alpha^{\prime}_{4}\\ \beta^{\prime}_{1}&-\alpha^{\prime}_{1}&-\alpha^{\prime}_{1}&-\alpha^{\prime}_{2}\end{array}\right)=gA(g^{-1})^{\otimes 2}\ \mbox{then}$
$\alpha^{\prime}_{1}=\frac{1}{\Delta}\left(-\beta_{1}\eta_{1}\xi_{1}^{2}+\alpha_{1}\eta_{2}\xi_{1}^{2}+2\alpha_{1}\eta_{1}\xi_{1}\xi_{2}+2\alpha_{2}\eta_{2}\xi_{1}\xi_{2}+\alpha_{2}\eta_{1}\xi_{2}^{2}+\alpha_{4}\eta_{2}\xi_{2}^{2}\right),$
$\alpha^{\prime}_{2}=\frac{-1}{\Delta}\left(\beta_{1}\eta_{1}^{2}\xi_{1}-2\alpha_{1}\eta_{1}\eta_{2}\xi_{1}-\alpha_{2}\eta_{2}^{2}\xi_{1}-\alpha_{1}\eta_{1}^{2}\xi_{2}-2\alpha_{2}\eta_{1}\eta_{2}\xi_{2}-\alpha_{4}\eta_{2}^{2}\xi_{2}\right),$
$\alpha^{\prime}_{4}=\frac{-1}{\Delta}\left(\beta_{1}\eta_{1}^{3}-3\alpha_{1}\eta_{1}^{2}\eta_{2}-3\alpha_{2}\eta_{1}\eta_{2}^{2}-\alpha_{4}\eta_{2}^{3}\right),$
$\beta^{\prime}_{1}=\frac{1}{\Delta}\left(\beta_{1}\xi_{1}^{3}-3\alpha_{1}\xi_{1}^{2}\xi_{2}-3\alpha_{2}\xi_{1}\xi_{2}^{2}-\alpha_{4}\xi_{2}^{3}\right).$
Case 1: $\alpha_{4}\neq 0$. It follows that if $\xi_{1}=0$ and $\eta_{2}=-\frac{\alpha_{2}}{\alpha_{4}}$, then $\alpha^{\prime}_{1}=0$. Therefore it is sufficient to consider the case $\alpha_{1}=0$.
Case 1-1: $\alpha_{2}\neq 0$. If $\xi_{2}=\eta_{1}=0$, then $\alpha^{\prime}_{1}=0,\ \alpha^{\prime}_{2}=\alpha_{2}\eta_{2},\ \alpha^{\prime}_{4}=\alpha_{4}\frac{\eta_{2}^{2}}{\xi_{1}},\ \beta^{\prime}_{1}=\beta_{1}\frac{\xi_{1}^{2}}{\eta_{2}}$. So one can make $\alpha^{\prime}_{2}=1,\ \alpha^{\prime}_{4}=1$. Therefore, without loss of generality, one can assume that $\alpha_{1}=0,\alpha_{2}=1,\alpha_{4}=1$ and the above system becomes
$\alpha^{\prime}_{1}=\frac{1}{\Delta}\left(-\beta_{1}\eta_{1}\xi_{1}^{2}+2\eta_{2}\xi_{1}\xi_{2}+\eta_{1}\xi_{2}^{2}+\eta_{2}\xi_{2}^{2}\right),$
$\alpha^{\prime}_{2}=\frac{-1}{\Delta}\left(\beta_{1}\eta_{1}^{2}\xi_{1}-\eta_{2}^{2}\xi_{1}-2\eta_{1}\eta_{2}\xi_{2}-\eta_{2}^{2}\xi_{2}\right),$
$\alpha^{\prime}_{4}=\frac{-1}{\Delta}\left(\beta_{1}\eta_{1}^{3}-3\eta_{1}\eta_{2}^{2}-\eta_{2}^{3}\right),$
$\beta^{\prime}_{1}=\frac{1}{\Delta}\left(\beta_{1}\xi_{1}^{3}-3\xi_{1}\xi_{2}^{2}-\xi_{2}^{3}\right).$
In the $\xi_{2}\eta_{1}=0$ case this yields the algebra $\left(\begin{array}[]{cccc}0&1&1&1\\ \beta_{1}&0&0&-1\end{array}\right)\simeq\left(\begin{array}[]{cccc}0&1&1&1\\ 4-\beta_{1}&0&0&-1\end{array}\right)$. Indeed, if $\eta_{1}=0$, then $\Delta=\xi_{1}\eta_{2}$ and
$\alpha^{\prime}_{1}=\xi_{1}\left(2\xi_{2}/\xi_{1}+(\xi_{2}/\xi_{1})^{2}\right)$ implies $\xi_{2}=-2\xi_{1},$
$\alpha^{\prime}_{2}=\eta_{2}\left(1+\xi_{2}/\xi_{1}\right)=-\eta_{2}\xi_{1}=1,$
$\alpha^{\prime}_{4}=\frac{1}{\xi_{1}}\left(\eta_{2}^{2}\right)=\frac{1}{\xi_{1}^{3}}=1.$ So, if $\xi_{1}=a,\xi_{2}=-2a$ and $\eta_{2}=-1/a$, where $a^{3}=1$, then
$\beta^{\prime}_{1}=\frac{1}{-1}(\beta_{1}a^{3}-12a^{3}+8a^{3})=4-\beta_{1}.$
In the $\xi_{2}=0$ case no new algebra appears.
In the $\xi_{2}\eta_{1}\neq 0$ case $\alpha^{\prime}_{1}=0$ is equivalent to $\frac{\eta_{2}}{\eta_{1}}(2\frac{\xi_{1}}{\xi_{2}}+1)-\beta_{1}(\frac{\xi_{1}}{\xi_{2}})^{2}+1=0$. If $2\frac{\xi_{1}}{\xi_{2}}+1=0$, then this yields $\left(\begin{array}[]{cccc}0&1&1&1\\ 4&0&0&-1\end{array}\right)\simeq\left(\begin{array}[]{cccc}0&1&1&1\\ 0&0&0&-1\end{array}\right)$.
If $2\frac{\xi_{1}}{\xi_{2}}+1\neq 0$, then $\frac{\eta_{2}}{\eta_{1}}=\frac{\beta_{1}(\frac{\xi_{1}}{\xi_{2}})^{2}-1}{2\frac{\xi_{1}}{\xi_{2}}+1}=\frac{\beta_{1}\xi^{2}-1}{2\xi+1}$, where $\xi=\frac{\xi_{1}}{\xi_{2}}$, and
$\alpha^{\prime}_{2}=\frac{\eta_{1}^{2}\xi_{2}}{\Delta}\left(-\beta_{1}\frac{\xi_{1}}{\xi_{2}}+2\frac{\eta_{2}}{\eta_{1}}+(1+\frac{\xi_{1}}{\xi_{2}})(\frac{\eta_{2}}{\eta_{1}})^{2}\right)=$
$\frac{\eta_{1}^{2}\xi_{2}}{\Delta}\left(-\beta_{1}\xi+2\frac{\beta_{1}\xi^{2}-1}{2\xi+1}+(1+\xi)(\frac{\beta_{1}\xi^{2}-1}{2\alpha_{2}\xi+1})^{2}\right)=$
$\frac{\eta_{1}^{2}\xi_{2}}{\Delta}\frac{\beta_{1}^{2}\xi^{5}+\beta_{1}^{2}\xi^{4}-2\beta_{1}\xi^{3}-4\beta_{1}\xi^{2}-(3+\beta_{1})\xi-1}{(2\xi+1)^{2}}=$
$\frac{\eta_{1}^{2}\xi_{2}}{\Delta}\frac{(\beta_{1}\xi^{3}-3\xi-1)(\beta_{1}\xi^{2}+\beta_{1}\xi+1)}{(2\xi+1)^{2}}.$
$\beta^{\prime}_{1}=\frac{\xi_{2}^{3}}{\Delta}(\beta_{1}\xi^{3}-3\xi-1)$
$\alpha^{\prime}_{4}=-\frac{\eta_{1}^{3}}{\Delta}(\beta_{1}-3(\frac{\eta_{2}}{\eta_{1}})^{2}-(\frac{\eta_{2}}{\eta_{1}})^{3})=\frac{\eta_{1}^{3}}{(2\xi+1)^{3}\Delta}(\beta_{1}^{3}\xi^{6}+6\beta_{1}^{2}\xi^{5}-20\beta_{1}\xi^{3}-15\beta_{1}\xi^{2}+6(1-\beta_{1})\xi+2-\beta_{1})$
Note that the equalities
$\Delta=\xi_{2}\eta_{1}(\xi\frac{\beta_{1}\xi^{2}-1}{2\xi+1}-1)=\xi_{2}\eta_{1}\frac{\beta_{1}\xi^{3}-3\xi-1}{2\xi+1},$
$P(t)=\beta_{1}^{3}t^{6}+6\beta_{1}^{2}t^{5}-20\beta_{1}t^{3}-15\beta_{1}t^{2}+6(1-\beta_{1})t+2-\beta_{1}=(\beta_{1}t^{3}-3t-1)(\beta_{1}^{2}t^{3}+6\beta_{1}t^{2}+3\beta_{1}t+\beta_{1}-2)$
are valid.
To make $\alpha^{\prime}_{2}=\alpha^{\prime}_{4}=1$ there should exist $\xi_{0}\neq-1/2$ such that $P_{1}(\xi_{0})P_{2}(\xi_{0})P_{3}(\xi_{0})\neq 0$, where $P_{1}(t)=\beta_{1}t^{2}+\beta_{1}t+1,P_{2}(t)=\beta_{1}t^{3}-3t-1,P_{3}(t)=\beta_{1}^{2}t^{3}+6\beta_{1}t^{2}+3\beta_{1}t+\beta_{1}-2$.
In this case
$\alpha^{\prime}_{2}=\frac{\eta_{1}^{2}\xi_{2}}{\Delta}\frac{(\beta_{1}\xi_{0}^{3}-3\xi_{0}-1)(\beta_{1}\xi_{0}^{2}+\beta_{1}\xi_{0}+1)}{(2\xi_{0}+1)^{2}}=\eta_{1}\frac{\beta_{1}\xi_{0}^{2}+\beta_{1}\xi_{0}+1}{2\xi_{0}+1}.$
Therefore if $\eta_{1}=\frac{2\xi_{0}+1}{\beta_{1}\xi_{0}^{2}+\beta_{1}\xi_{0}+1}$, then $\alpha^{\prime}_{2}=1$.
Equality $\alpha^{\prime}_{4}=\frac{\eta_{1}^{3}}{(2\xi_{0}+1)^{3}\Delta}P(\xi_{0})=1$ is equivalent to $\xi_{2}=\frac{\beta_{1}^{2}\xi_{0}^{3}+6\beta_{1}\xi_{0}^{2}+3\beta_{1}\xi_{0}+\beta_{1}-2}{(\beta_{1}\xi_{0}^{2}+\beta_{1}\xi_{0}+1)^{2}}$.
In this case
$\beta^{\prime}_{1}=\frac{\xi_{2}^{3}}{\Delta}(\beta_{1}\xi_{0}^{3}-3\xi_{0}-1)=\frac{(2\xi_{0}+1)\xi_{2}^{2}}{\eta_{1}}=\frac{(\beta_{1}^{2}\xi_{0}^{3}+6\beta_{1}\xi_{0}^{2}+3\beta_{1}\xi_{0}+\beta_{1}-2)^{2}}{(\beta_{1}\xi_{0}^{2}+\beta_{1}\xi_{0}+1)^{3}}$
and we obtain MSC
$\left(\begin{array}[]{cccc}0&1&1&1\\ 4-\beta_{1}&0&0&-1\end{array}\right)\simeq A_{10}(\mathbf{c})=\left(\begin{array}[]{cccc}0&1&1&1\\ \beta_{1}&0&0&-1\end{array}\right)\simeq\left(\begin{array}[]{cccc}0&1&1&1\\ \beta^{\prime}_{1}(a)&0&0&-1\end{array}\right),$
where $\frac{-1}{2}\neq a\in\mathbb{F}$, $(\beta_{1}a^{3}-3a-1)(\beta_{1}a^{2}+\beta_{1}a+1)(\beta_{1}^{2}a^{3}+6\beta_{1}a^{2}+3\beta_{1}a+\beta_{1}-2)\neq 0$, and
$\beta^{\prime}_{1}(t)=\frac{(\beta_{1}^{2}t^{3}+6\beta_{1}t^{2}+3\beta_{1}t+\beta_{1}-2)^{2}}{(\beta_{1}t^{2}+\beta_{1}t+1)^{3}}.$
It follows that $\beta^{\prime}_{1}(\frac{-1}{2})=4-\beta_{1}$.
Case 1-2: $\alpha_{2}=0$. The system becomes
$\begin{array}[]{ll}\alpha^{\prime}_{1}=\frac{1}{\Delta}\left(-\beta_{1}\eta_{1}\xi_{1}^{2}+\alpha_{4}\eta_{2}\xi_{2}^{2}\right),&\alpha^{\prime}_{2}=\frac{-1}{\Delta}\left(\beta_{1}\eta_{1}^{2}\xi_{1}-\alpha_{4}\eta_{2}^{2}\xi_{2}\right),\\ \alpha^{\prime}_{4}=\frac{-1}{\Delta}\left(\beta_{1}\eta_{1}^{3}-\alpha_{4}\eta_{2}^{3}\right),&\beta^{\prime}_{1}=\frac{1}{\Delta}\left(\beta_{1}\xi_{1}^{3}-\alpha_{4}\xi_{2}^{3}\right)\end{array}.$
To make $\alpha^{\prime}_{1}=\alpha^{\prime}_{2}=0$ we must have $\xi_{1}=\eta_{2}=0$ or $\xi_{2}=\eta_{1}=0$. If $\xi_{2}=\eta_{1}=0$ and $\xi_{1}=\alpha_{4}\eta_{2}^{2}$, then $\alpha^{\prime}_{4}=\frac{-1}{\Delta}\left(\beta_{1}\eta_{1}^{3}-\alpha_{4}\eta_{2}^{3}\right)=\frac{-1}{\xi_{1}}(-\alpha_{4}\eta_{2}^{2})=1$ and $\beta^{\prime}_{1}=\frac{1}{\Delta}\left(\beta_{1}\xi_{1}^{3}-\alpha_{4}\xi_{2}^{3}\right)=\frac{\xi_{1}^{2}}{\eta_{2}}\beta_{1}=\beta_{1}\alpha_{4}^{2}\eta_{2}^{3}$, so it yields the algebra
$A_{11}(\mathbf{c})=\left(\begin{array}[]{cccc}0&0&0&1\\ \beta_{1}&0&0&0\end{array}\right)\simeq\left(\begin{array}[]{cccc}0&0&0&1\\ a^{3}\beta_{1}&0&0&0\end{array}\right),$
where $\mathbf{c}=\beta_{1}\in\mathbb{F}$ and $0\neq a\in\mathbb{F}$.
Case 2: $\alpha_{4}=0$. The system becomes
$\alpha^{\prime}_{1}=\frac{1}{\Delta}\left(-\beta_{1}\eta_{1}\xi_{1}^{2}+\alpha_{1}\eta_{2}\xi_{1}^{2}+2\alpha_{1}\eta_{1}\xi_{1}\xi_{2}+2\alpha_{2}\eta_{2}\xi_{1}\xi_{2}+\alpha_{2}\eta_{1}\xi_{2}^{2}\right),$
$\alpha^{\prime}_{2}=\frac{-1}{\Delta}\left(\beta_{1}\eta_{1}^{2}\xi_{1}-2\alpha_{1}\eta_{1}\eta_{2}\xi_{1}-\alpha_{2}\eta_{2}^{2}\xi_{1}-\alpha_{1}\eta_{1}^{2}\xi_{2}-2\alpha_{2}\eta_{1}\eta_{2}\xi_{2}\right),$
$\alpha^{\prime}_{4}=\frac{-\eta_{1}}{\Delta}\left(\beta_{1}\eta_{1}^{2}-3\alpha_{1}\eta_{1}\eta_{2}-3\alpha_{2}\eta_{2}^{2}\right),$
$\beta^{\prime}_{1}=\frac{\xi_{1}}{\Delta}\left(\beta_{1}\xi_{1}^{2}-3\alpha_{1}\xi_{1}\xi_{2}-3\alpha_{2}\xi_{2}^{2}\right).$
If $\eta_{1}=0$, then $\Delta=\xi_{1}\eta_{2}$, $\alpha^{\prime}_{4}=0$ and
$\alpha^{\prime}_{1}=\xi_{1}\left(\alpha_{1}+2\alpha_{2}\frac{\xi_{2}}{\xi_{1}}\right),$
$\alpha^{\prime}_{2}=\alpha_{2}\eta_{2},$
$\beta^{\prime}_{1}=\frac{\xi_{1}^{2}}{\eta_{2}}\left(\beta_{1}-3\alpha_{1}\frac{\xi_{2}}{\xi_{1}}-3\alpha_{2}(\frac{\xi_{2}}{\xi_{1}})^{2}\right).$
Case 2-1: $\alpha_{2}\neq 0$. If $\frac{\xi_{2}}{\xi_{1}}=\frac{-\alpha_{1}}{2\alpha_{2}}$, then $\alpha^{\prime}_{1}=0$ $\alpha^{\prime}_{2}\neq 0$. After $\alpha^{\prime}_{2}$ can be made 11 to get
$A_{12}(\mathbf{c})=\left(\begin{array}[]{cccc}0&1&1&0\\ \beta_{1}&0&0&-1\end{array}\right)\simeq\left(\begin{array}[]{cccc}0&1&1&0\\ a^{2}\beta_{1}&0&0&-1\end{array}\right),$ where $\mathbf{c}=\beta_{1}\in\mathbb{F}$ and $0\neq a\in\mathbb{F}$.
Case 2-2: $\alpha_{2}=0$. The corresponding system is
$\alpha^{\prime}_{1}=\frac{\xi_{1}}{\Delta}\left(-\beta_{1}\eta_{1}\xi_{1}+\alpha_{1}\eta_{2}\xi_{1}+2\alpha_{1}\eta_{1}\xi_{2}\right),$
$\alpha^{\prime}_{2}=\frac{-\eta_{1}}{\Delta}\left(\beta_{1}\eta_{1}\xi_{1}-2\alpha_{1}\eta_{2}\xi_{1}-\alpha_{1}\eta_{1}\xi_{2}\right),$
$\alpha^{\prime}_{4}=\frac{-\eta_{1}^{2}}{\Delta}\left(\beta_{1}\eta_{1}-3\alpha_{1}\eta_{2}\right),$
$\beta^{\prime}_{1}=\frac{\xi_{1}^{2}}{\Delta}\left(\beta_{1}\xi_{1}-3\alpha_{1}\xi_{2}\right).$
If $\eta_{1}=0$, then $\alpha^{\prime}_{2}=\alpha^{\prime}_{4}=0$ and $\alpha^{\prime}_{1}=\xi_{1}\alpha_{1},\beta^{\prime}_{1}=\frac{\xi^{2}_{1}}{\eta_{2}}\left(\beta_{1}-3\alpha_{1}\frac{\xi_{2}}{\xi_{1}}\right).$ Therefore, if $\alpha_{1}\neq 0$, we may assume $\alpha^{\prime}_{1}=1,\beta^{\prime}_{1}=0$ to get $A^{\prime}=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&-1&-1&0\end{array}\right)$, which is isomorphic to $A_{12}(0),$ if $\alpha_{1}=0$, then $\alpha^{\prime}_{1}=0$ and we may assume $\beta^{\prime}_{1}=1$ to get $A=\left(\begin{array}[]{cccc}0&0&0&0\\ 1&0&0&0\end{array}\right),$ which is isomorphic to $A_{11}(0).$
We now consider the case $char(\mathbb{F})=2$. The approach is the same.
The fifth subset in the $char(\mathbb{F})=2$ case.
The same approach applies.
Case 1: $\alpha_{4}\neq 0$. If $\xi_{1}=0$ and $\eta_{2}=-\frac{\alpha_{2}}{\alpha_{4}}$, then $\alpha^{\prime}_{1}=0$. So further we consider only $\alpha_{1}=0$ case.
Case 1.1: $\alpha_{2}\neq 0$. If $\xi_{2}=\eta_{1}=0$, then $\alpha^{\prime}_{1}=0,\ \alpha^{\prime}_{2}=\alpha_{2}\eta_{2},\ \alpha^{\prime}_{4}=\alpha_{4}\frac{\eta_{2}^{2}}{\xi_{1}},\ \beta^{\prime}_{1}=\beta_{1}\frac{\xi_{1}^{2}}{\eta_{2}}$. Therefore, without loss of generality, one can assume that $\alpha_{1}=0,\ \alpha_{2}=1,\ \alpha_{4}=1$. The corresponding system is:
$\alpha^{\prime}_{1}=\frac{1}{\Delta}\left(-\beta_{1}\eta_{1}\xi_{1}^{2}+\eta_{1}\xi_{2}^{2}+\eta_{2}\xi_{2}^{2}\right),$
$\alpha^{\prime}_{2}=\frac{-1}{\Delta}\left(\beta_{1}\eta_{1}^{2}\xi_{1}-\eta_{2}^{2}\xi_{1}-\eta_{2}^{2}\xi_{2}\right),$
$\alpha^{\prime}_{4}=\frac{-1}{\Delta}\left(\beta_{1}\eta_{1}^{3}-\eta_{1}\eta_{2}^{2}-\eta_{2}^{3}\right),$
$\beta^{\prime}_{1}=\frac{1}{\Delta}\left(\beta_{1}\xi_{1}^{3}-\xi_{1}\xi_{2}^{2}-\xi_{2}^{3}\right).$
If $\xi_{2}=\eta_{1}=0$, then $\Delta=\xi_{1}\eta_{2}$ and $\alpha^{\prime}_{1}=0,$ $\alpha^{\prime}_{2}=\eta_{2}\xi_{1},$ $\alpha^{\prime}_{4}=\frac{\eta_{2}^{2}}{\xi_{1}}$, $\beta^{\prime}_{1}=\frac{\xi_{1}^{2}}{\eta_{2}}\beta_{1}.$ By $\eta_{2}=1,\xi_{1}=1$ we obtain $\left(\begin{array}[]{cccc}0&1&1&1\\ \beta_{1}&0&0&-1\end{array}\right)\simeq\left(\begin{array}[]{cccc}0&1&1&1\\ \beta_{1}&0&0&-1\end{array}\right)$.
In the $\xi_{2}\eta_{1}\neq 0$ case $\alpha^{\prime}_{1}=0$ if and only if $\frac{\eta_{2}}{\eta_{1}}=\beta_{1}\xi^{2}-1$, where $\xi=\frac{\xi_{1}}{\xi_{2}}$. In this case
$\alpha^{\prime}_{2}=\frac{\eta_{1}^{2}\xi_{2}}{\Delta}\left(-\beta_{1}\frac{\xi_{1}}{\xi_{2}}+(1+\frac{\xi_{1}}{\xi_{2}})(\frac{\eta_{2}}{\eta_{1}})^{2}\right)=$
$\frac{\eta_{1}^{2}\xi_{2}}{\Delta}\left(-\beta_{1}\xi+(1+\xi)(\beta_{1}\xi^{2}-1)^{2}\right)=$
$\frac{\eta_{1}^{2}\xi_{2}}{\Delta}(\beta_{1}^{2}\xi^{5}+\beta_{1}^{2}\xi^{4}-(3+\beta_{1})\xi-1)=$
$\frac{\eta_{1}^{2}\xi_{2}}{\Delta}(\beta_{1}\xi^{3}-\xi-1)(\beta_{1}\xi^{2}+\beta_{1}\xi+1).$
$\beta^{\prime}_{1}=\frac{\xi_{2}^{3}}{\Delta}(\beta_{1}\xi^{3}-\xi-1)$
$\alpha^{\prime}_{4}=-\frac{\eta_{1}^{3}}{\Delta}(\beta_{1}-(\frac{\eta_{2}}{\eta_{1}})^{2}-(\frac{\eta_{2}}{\eta_{1}})^{3})=\frac{\eta_{1}^{3}}{\Delta}(\beta_{1}^{3}\xi^{6}-\beta_{1}\xi^{2}-\beta_{1})$
Note that the equalities
$\Delta=\xi_{2}\eta_{1}(\xi(\beta_{1}\xi^{2}-1)-1)=\xi_{2}\eta_{1}(\beta_{1}\xi^{3}-\xi-1),$
$P(t)=\beta_{1}^{3}t^{6}-\beta_{1}t^{2}-\beta_{1}=(\beta_{1}t^{3}+t+1)^{2}\beta_{1}$
are valid.
To make $\alpha^{\prime}_{2}=\alpha^{\prime}_{4}=1$ one should have $\beta_{1}\neq 0$ and there should exist $\xi_{0}$ such that $P_{1}(\xi_{0})P_{2}(\xi_{0})\neq 0$, where $P_{1}(t)=\beta_{1}t^{2}+\beta_{1}t+1,P_{2}(t)=\beta_{1}t^{3}+t+1$.
In this case
$\alpha^{\prime}_{2}=\frac{\eta_{1}^{2}\xi_{2}}{\Delta}(\beta_{1}\xi_{0}^{3}-\xi_{0}-1)(\beta_{1}\xi_{0}^{2}+\beta_{1}\xi_{0}+1)=\eta_{1}(\beta_{1}\xi_{0}^{2}+\beta_{1}\xi_{0}+1).$
Therefore, if $\eta_{1}=\frac{1}{\beta_{1}\xi_{0}^{2}+\beta_{1}\xi_{0}+1}$, then $\alpha^{\prime}_{2}=1$.
Equality $\alpha^{\prime}_{4}=\frac{\eta_{1}^{3}}{(2\xi_{0}+1)^{3}\Delta}P(\xi_{0})=1$ is equivalent to $\xi_{2}=\beta_{1}\frac{\beta_{1}\xi_{0}^{3}+\xi_{0}+1}{(\beta_{1}\xi_{0}^{2}+\beta_{1}\xi_{0}+1)^{2}}$.
In this case
$\beta^{\prime}_{1}=\frac{\xi_{2}^{3}}{\Delta}(\beta_{1}\xi_{0}^{3}-\xi_{0}-1)=\beta_{1}^{2}\frac{(\beta_{1}\xi_{0}^{3}+\xi_{0}+1)^{2}}{(\beta_{1}\xi_{0}^{2}+\beta_{1}\xi_{0}+1)^{3}}$
and this yields MSC
$A_{8.2}(\mathbf{c})=\left(\begin{array}[]{cccc}0&1&1&1\\ \beta_{1}&0&0&-1\end{array}\right)\simeq\left(\begin{array}[]{cccc}0&1&1&1\\ \beta^{\prime}_{1}(a)&0&0&-1\end{array}\right),$
where $(\beta_{1}a^{3}+a+1)(\beta_{1}a^{2}+\beta_{1}a+1)\beta_{1}\neq 0$ and
$\beta^{\prime}_{1}(t)=\beta_{1}^{2}\frac{(\beta_{1}t^{3}+t+1)^{2}}{(\beta_{1}t^{2}+\beta_{1}t+1)^{3}}.$
Case 1.2: $\alpha_{2}=0$. The system becomes
$\begin{array}[]{ll}\alpha^{\prime}_{1}=\frac{1}{\Delta}\left(-\beta_{1}\eta_{1}\xi_{1}^{2}+\alpha_{4}\eta_{2}\xi_{2}^{2}\right),&\alpha^{\prime}_{2}=\frac{-1}{\Delta}\left(\beta_{1}\eta_{1}^{2}\xi_{1}-\alpha_{4}\eta_{2}^{2}\xi_{2}\right),\\ \alpha^{\prime}_{4}=\frac{-1}{\Delta}\left(\beta_{1}\eta_{1}^{3}-\alpha_{4}\eta_{2}^{3}\right),&\beta^{\prime}_{1}=\frac{1}{\Delta}\left(\beta_{1}\xi_{1}^{3}-\alpha_{4}\xi_{2}^{3}\right)\end{array}$
If $\xi_{2}=\eta_{1}=0$, then $\alpha^{\prime}_{4}=\frac{-1}{\Delta}\left(\beta_{1}\eta_{1}^{3}-\alpha_{4}\eta_{2}^{3}\right)=\frac{\alpha_{4}\eta_{2}^{2}}{\xi_{1}}$ and $\beta^{\prime}_{1}=\frac{1}{\Delta}\left(\beta_{1}\xi_{1}^{3}-\alpha_{4}\xi_{2}^{3}\right)=\frac{\xi_{1}^{2}}{\eta_{2}}\beta_{1}=\beta_{1}\alpha_{4}^{2}\eta_{2}^{3}$. Therefore one has
$A_{9.2}(\mathbf{c})=\left(\begin{array}[]{cccc}0&0&0&1\\ \beta_{1}&0&0&0\end{array}\right)\simeq\left(\begin{array}[]{cccc}0&0&0&1\\ a^{3}\beta_{1}&0&0&0\end{array}\right),$
where $\mathbf{c}=\beta_{1}\in\mathbb{F}$ and $0\neq a\in\mathbb{F}$.
Case 2: $\alpha_{4}=0$. The system becomes
$\alpha^{\prime}_{1}=\frac{1}{\Delta}\left(-\beta_{1}\eta_{1}\xi_{1}^{2}+\alpha_{1}\eta_{2}\xi_{1}^{2}+\alpha_{2}\eta_{1}\xi_{2}^{2}\right),$
$\alpha^{\prime}_{2}=\frac{-1}{\Delta}\left(\beta_{1}\eta_{1}^{2}\xi_{1}-\alpha_{2}\eta_{2}^{2}\xi_{1}-\alpha_{1}\eta_{1}^{2}\xi_{2}\right),$
$\alpha^{\prime}_{4}=\frac{-\eta_{1}}{\Delta}\left(\beta_{1}\eta_{1}^{2}-\alpha_{1}\eta_{1}\eta_{2}-\alpha_{2}\eta_{2}^{2}\right),$
$\beta^{\prime}_{1}=\frac{\xi_{1}}{\Delta}\left(\beta_{1}\xi_{1}^{2}-\alpha_{1}\xi_{1}\xi_{2}-\alpha_{2}\xi_{2}^{2}\right).$
If $\eta_{1}=0$, then $\alpha^{\prime}_{4}=0$, $\Delta=\xi_{1}\eta_{2}$ and
$\alpha^{\prime}_{1}=\xi_{1}\alpha_{1},$
$\alpha^{\prime}_{2}=\alpha_{2}\eta_{2},$
$\beta^{\prime}_{1}=\frac{\xi_{1}^{2}}{\eta_{2}}\left(\beta_{1}-\alpha_{1}\frac{\xi_{2}}{\xi_{1}}-\alpha_{2}(\frac{\xi_{2}}{\xi_{1}})^{2}\right).$
Case 2-1: $\alpha_{1}\neq 0,\alpha_{2}\neq 0$. We may assume $\alpha^{\prime}_{1}=1$ $\alpha^{\prime}_{2}=1$ to get
$A_{10,2}(\mathbf{c})=\left(\begin{array}[]{cccc}1&1&1&0\\ \beta_{1}&1&1&1\end{array}\right)\simeq\left(\begin{array}[]{cccc}1&1&1&0\\ a^{2}+a+\beta_{1}&1&1&1\end{array}\right),$ where $\mathbf{c}=\beta_{1}\in\mathbb{F}$ and $a\in\mathbb{F}$.
Case 2-2: $\alpha_{1}=0,\alpha_{2}\neq 0$. In this we obtain
$A_{11,2}(\mathbf{c})=\left(\begin{array}[]{cccc}0&1&1&0\\ \beta_{1}&0&0&1\end{array}\right)\simeq\left(\begin{array}[]{cccc}0&1&1&0\\ b^{2}(\beta_{1}+a^{2})&0&0&1\end{array}\right),$ where $\mathbf{c}=\beta_{1}\in\mathbb{F}$ and $a,b\in\mathbb{F},b\neq 0$.
Case 2-3: $\alpha_{1}\neq 0,\alpha_{2}=0$. This yields
$A^{\prime}=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&-1&-1&0\end{array}\right)$, which is isomorphic to $A_{11,2}(0)$.
Case 2-4: $\alpha_{1}=0,\alpha_{2}=0$. In this case one has $A=\left(\begin{array}[]{cccc}0&0&0&0\\ 1&0&0&0\end{array}\right)\simeq A_{9,2}(0).$
The fifth subset in the $char(\mathbb{F})=3$ case.
The proof is similar to that of $char(\mathbb{F})\neq 2,3$ except for the following case.
Case 2-2: $\alpha_{2}=0$. The corresponding system is
$\alpha^{\prime}_{1}=\frac{\xi_{1}}{\Delta}\left(-\beta_{1}\eta_{1}\xi_{1}+\alpha_{1}\eta_{2}\xi_{1}+2\alpha_{1}\eta_{1}\xi_{2}\right),$
$\alpha^{\prime}_{2}=\frac{-\eta_{1}}{\Delta}\left(\beta_{1}\eta_{1}\xi_{1}-2\alpha_{1}\eta_{2}\xi_{1}-\alpha_{1}\eta_{1}\xi_{2}\right),$
$\alpha^{\prime}_{4}=\frac{-\eta_{1}^{2}}{\Delta}\left(\beta_{1}\eta_{1}-3\alpha_{1}\eta_{2}\right),$
$\beta^{\prime}_{1}=\frac{\xi_{1}^{2}}{\Delta}\left(\beta_{1}\xi_{1}-3\alpha_{1}\xi_{2}\right).$
If $\eta_{1}=0$, then $\alpha^{\prime}_{2}=\alpha^{\prime}_{4}=0$ and $\alpha^{\prime}_{1}=\xi_{1}\alpha_{1},\beta^{\prime}_{1}=\frac{\xi^{2}_{1}}{\eta_{2}}\beta_{1}.$ Therefore, one can have $A^{\prime}=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&-1&-1&0\end{array}\right)\simeq A_{11,3}(0)$, or $A_{12,3}=\left(\begin{array}[]{cccc}1&0&0&0\\ 1&-1&-1&0\end{array}\right)\simeq A_{9,3}(0)$ or $A=\left(\begin{array}[]{cccc}0&0&0&0\\ 1&0&0&0\end{array}\right)\simeq A_{10,3}(0)$.
Conclusion. In the Theorem 1.1 the items $A_{13},A_{12,2},A_{12,3},A_{13,3}$ should be omitted and the items $A_{10},A_{11}$, $A_{8,2},A_{9,2}$, $A_{10,3},A_{11,3}$ should be replaced by the following items, respectively:
  • $A_{10}(\mathbf{c})=\left(\begin{array}[]{cccc}0&1&1&1\\ \beta_{1}&0&0&-1\end{array}\right)\simeq\left(\begin{array}[]{cccc}0&1&1&1\\ \beta^{\prime}_{1}(a)&0&0&-1\end{array}\right),$ where
    $(\beta_{1}a^{3}-3a-1)(\beta_{1}a^{2}+\beta_{1}a+1)(\beta_{1}^{2}a^{3}+6\beta_{1}a^{2}+3\beta_{1}a+\beta_{1}-2)\neq 0$, $\beta^{\prime}_{1}(t)=\frac{(\beta_{1}^{2}t^{3}+6\beta_{1}t^{2}+3\beta_{1}t+\beta_{1}-2)^{2}}{(\beta_{1}t^{2}+\beta_{1}t+1)^{3}}$ and $\beta^{\prime}(\frac{-1}{2})$ is considered to be $4-\beta_{1}$.
  • $A_{11}(\mathbf{c})=\left(\begin{array}[]{cccc}0&0&0&1\\ \beta_{1}&0&0&0\end{array}\right)\simeq\left(\begin{array}[]{cccc}0&0&0&1\\ a^{3}\beta_{1}&0&0&0\end{array}\right),$ where $\mathbf{c}=\beta_{1}$ and $0\neq a\in\mathbb{F}$.
  • $A_{8,2}(\mathbf{c})=\left(\begin{array}[]{cccc}0&1&1&1\\ \beta_{1}&0&0&-1\end{array}\right)\simeq\left(\begin{array}[]{cccc}0&1&1&1\\ \beta^{\prime}_{1}(a)&0&0&-1\end{array}\right),$ where
    $(\beta_{1}a^{3}+a+1)(\beta_{1}a^{2}+\beta_{1}a+1)\beta_{1}\neq 0$, and $\beta^{\prime}_{1}(t)=\beta_{1}^{2}\frac{(\beta_{1}t^{3}+t+1)^{2}}{(\beta_{1}t^{2}+\beta_{1}t+1)^{3}}.$
  • $A_{9,2}(\mathbf{c})=\left(\begin{array}[]{cccc}0&0&0&1\\ \beta_{1}&0&0&0\end{array}\right)\simeq\left(\begin{array}[]{cccc}0&0&0&1\\ a^{3}\beta_{1}&0&0&0\end{array}\right),$ where $\mathbf{c}=\beta_{1}\in\mathbb{F}$ and $0\neq a\in\mathbb{F}$.
  • $A_{9,3}(\mathbf{c})=\left(\begin{array}[]{cccc}0&1&1&1\\ \beta_{1}&0&0&-1\end{array}\right)\simeq\left(\begin{array}[]{cccc}0&1&1&1\\ \beta^{\prime}_{1}(a)&0&0&-1\end{array}\right),$ where
    $1\neq a\in\mathbb{F}$, $(\beta_{1}a^{3}-1)(\beta_{1}a^{2}+\beta_{1}a+1)(\beta_{1}^{2}a^{3}+\beta_{1}+1)\neq 0$, $\beta^{\prime}_{1}(t)=\frac{(\beta_{1}^{2}t^{3}+\beta_{1}+1)^{2}}{(\beta_{1}t^{2}+\beta_{1}t+1)^{3}}$ and $\beta^{\prime}(1)$ is considered to be $2\beta_{1}+1$.
  • $A_{10,3}(\mathbf{c})=\left(\begin{array}[]{cccc}0&0&0&1\\ \beta_{1}&0&0&0\end{array}\right)\simeq\left(\begin{array}[]{cccc}0&0&0&1\\ a^{3}\beta_{1}&0&0&0\end{array}\right),$ where $\mathbf{c}=\beta_{1}$ and $0\neq a\in\mathbb{F}$.
3. On the Number of Non-Isomorphic Two-Dimensional Algebras over a Finite Field
Let $|\mathbb{F}|=q=p^{n}$ and suppose that $\mathbb{F}^{*}=\langle s\rangle=\{1,s,\ldots,s^{q-1}\}$. Since $p$ and q1q-1 are coprime, it follows that $\mathbb{F}^{*}=\langle s^{p}\rangle$ as well.
We denote by $|A_{i}|$ the number of non-isomorphic algebras in the class $A_{i}$.
Proposition 3.1.
The following equalities hold::
  • $|A_{1}|=q^{4},|A_{2}|=q^{2}(q-1),|A_{3}|=3q^{2},|A_{4}|=q^{2},|A_{5}|=q,|A_{6}|=q(q-1),|A_{7}|=3q,|A_{8}|=q,|A_{9}|=1,|A_{11}|=\left\{\begin{array}[]{ll}2&,\,\,if\,\,3\,\,divides\,\,q-2,\\ 4&,\,\,otherwise.\end{array}\right.,|A_{12}|=3$ and the total number of them, including the trivial algebra, is
    $\left\{\begin{array}[]{ll}q^{4}+q^{3}+4q^{2}+4q+7&,\,\,if\,\,3\,\,divides\,\,q-2,\\ q^{4}+q^{3}+4q^{2}+4q+9&,\,\,otherwise.\end{array}\right$.
  • $|A_{1,2}|=q^{4},|A_{2,2}|=q^{2}(q-1)+q,|A_{3,2}|=2q^{2},|A_{4,2}|=2q^{2}-q,|A_{5,2}|=q(q-1)+1,|A_{6,2}|=2q,|A_{7,2}|=2q-1,|A_{9,2}|=\left\{\begin{array}[]{ll}2&,\,if\,q=2^{2k-1},\\ 4&,\,if\,q=2^{2k}\end{array}\right.,|A_{10,2}|=2,|A_{11,2}|=1$ and the total number of them, including the trivial algebra, is
    $\left\{\begin{array}[]{ll}q^{4}+q^{3}+4q^{2}+3q+6&,\,if\,q=2^{2k-1},\\ q^{4}+q^{3}+4q^{2}+3q+7&,\,if\,q=2^{2k}.\end{array}\right$.
  • $|A_{1,3}|=q^{4},|A_{2,3}|=q^{2}(q-1),|A_{3,3}|=3q^{2},|A_{4,3}|=q^{2},|A_{5,3}|=q,|A_{6,3}|=q(q-1),|A_{7,3}|=3q,|A_{8,3}|=q,|A_{10,3}|=2,|A_{11,3}|=3$ and the total number of them, including the trivial algebra, is
    $q^{4}+q^{3}+4q^{2}+4q+6.$
Proof.
Many of the above equalities can be verified directly. We discuss only some representative cases.
  • In the class $A_{3}$ In the class $A_{3}$, the parameter $a^{2}\alpha_{4}$ can be reduced to 0, 11, or $s$, while $\alpha_{1}$ and $\beta_{2}$ are free parameters. Therefore $|A_{3}|=3q^{2}.$ Similarly, $|A_{7}|=3q$. Since $1=s^{q-1}=ss^{q-2}=s^{2}s^{q-3}=s^{2}s^{2(q-2)}$ one concludes that if 33 divides q2q-2 then $|A_{11}|=2$, otherwise $|A_{11}|=4$. In $A_{12}$ case 22 does not divide q2q-2, therefore $|A_{12}|=3.$
  • Let $r_{a}=|R_{a}|$, where $R_{a}$ stands for the range the map $x\rightarrow x^{2}+ax$ from $\mathbb{F}$ to itself. It is clear that $R_{a}$ is closed with respect to the addition and $r_{0}=q$ and $r_{a}=q/2$ if $a\neq 0$. If $b$ is not in $R_{a}$, then $R_{a}$ and $b+R_{a}$ are disjoint subsets. Therefore, in $A_{4,2}$ case, if $\beta_{2}=1$, there are $q$ non-isomorphic algebras, for each $\beta_{2}\neq 1$ there are $2q$ non-isomorphic algebras, so the total number is $|A_{4,2}|=q+2(q-1)q=2q^{2}-q$. The situation in $A_{7,2}$ is similar that of $A_{4,2}$, so $|A_{7,2}|=1+2(q-1)=2q-1$.
    If $q=2^{2k-1}$, then 33 divides q2q-2 and due to $1=s(s^{\frac{q-2}{3}})^{3},s=s^{2}(s^{\frac{q-2}{3}})^{3}$ one has $|A_{9,2}|=2$, otherwise $|A_{9,2}|=4.$
    One has $|A_{10,2}|=2$ since $\beta_{1}$ can be reduced to either 0 or 11. In $A_{11,2}$ case $\beta_{1}$ can be made 0.
  • In $A_{10,3}$ case $|A_{10,3}|=2$ as far as $1=s^{q-1}=ss^{q-2}=ss^{2q-3}=s^{2}s^{q-3}$ and 33 divides q2q-2 and 2q32q-3. $|A_{11,3}|=3$ as far as 22 does not divide q2q-2.
Relaying on the above Lemma one can conclude:
Theorem 3.2.
Let $\mathbb{F}$ be a finite field with $|\mathbb{F}|=q$. The number of non-isomorphic two-dimensional algebras over $\mathbb{F}$ is given as follows:
  • $char(\mathbb{F})\neq 2,3$ case.
    $\left\{\begin{array}[]{ll}q^{4}+q^{3}+4q^{2}+4q+7+|A_{10}|&,\,\,if\,\,3\,\,divides\,\,q-2,\\ q^{4}+q^{3}+4q^{2}+4q+9+|A_{10}|&,\,\,otherwise.\end{array}\right$.
  • $char(\mathbb{F})=2$ case.
    $\left\{\begin{array}[]{ll}q^{4}+q^{3}+4q^{2}+3q+6+|A_{8,2}|&,\,if\,q=2^{2k-1},\\ q^{4}+q^{3}+4q^{2}+3q+7+|A_{8,2}|&,\,if\,q=2^{2k}.\end{array}\right$.
  • $char(\mathbb{F})=3$ case.
    $q^{4}+q^{3}+4q^{2}+4q+6+|A_{9,3}|.$
Remark $3.$
[…]
$f(f(a,s),t)=f(a,f(s,t)).$
Hence there is a need for an investigation of the orbit structure of this action.